What is the pH when 23.4 mL of 0.250 M Ca(OH)2 is diluted to 1.5 L?
1
Expert's answer
2020-05-15T10:23:35-0400

C1=n/V1=> n(Ca(OH)2)=C1*V1=0,25*0,0234=0,00585 moll

Ca(OH)2=> 2OH-+Ca2+

n(OH-)=0,00585*2=0,0117 moll

c(OH-)=0,0117/1,5=0,0078 moll/l

pOH=-lg(c(OH-))=2,11

pH=14-pOH=11,89